OpenStudy (anonymous):

If this calculation continued forever, what would you expect the answer to be: 1/3 + 1/9 + 1/27 + 1/81 + 1/243 + ....

OpenStudy (anonymous):

How can one do that?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I must of forgotten how its done

OpenStudy (anonymous):

I have one more question in relation to another math problem

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The power sign increases by one

OpenStudy (anonymous):

The n

OpenStudy (anonymous):

Does it have an answer because in the textbook the answer is 1/2

OpenStudy (anonymous):

What does the R stand for and the 'a' and 'k'?

OpenStudy (anonymous):

I sometimes skip class :)

terenzreignz (terenzreignz):

Speaking from experience, @precal ? :D

OpenStudy (anonymous):

I shouldn't that's why now I'm trying to focus more now on my studies

OpenStudy (anonymous):

Two degrees in what subject?

terenzreignz (terenzreignz):

Don't worry, @precal I'm a good (school) boy. @Chad123 Don't you think it's odd that you can actually sum up an infinite number of addends and you still get an answer? :D

OpenStudy (anonymous):

Maybe you did something else other than maths

terenzreignz (terenzreignz):

Yeah, there should be k = 1 ad the bottom of the big sigma :D

terenzreignz (terenzreignz):

if the exponent is n-1, it won't really matter :D

terenzreignz (terenzreignz):

Okay, now I see it :D

terenzreignz (terenzreignz):

I fix boredom by deriving :D \[\huge S = \frac13 + \frac19 + \frac1{27}+\frac1{81}+\frac1{243}...\]

terenzreignz (terenzreignz):

Here's an idea, @Chad123 Why don't we multiply both sides of the equation by 3? \[\huge 3S = 1+\frac13+\frac19+\frac1{27}+\frac1{81}+\frac1{243}...\]

terenzreignz (terenzreignz):

And then bring the 1 to the left side... \[\huge 3S - 1 =\frac13 + \frac19 + \frac1{27}+\frac1{81}+\frac1{243}... \]

terenzreignz (terenzreignz):

If you notice, the right-side is once again equal to S :D \[\huge 3S -1 = S\] Now it's just a linear equation in one variable. Get to it, champ, and solve for S :)

terenzreignz (terenzreignz):

Because we started with this\[\huge S = \frac13 + \frac19 + \frac1{27}+\frac1{81}+\frac1{243}...\]

terenzreignz (terenzreignz):

Eureka :)

OpenStudy (anonymous):

Is there another way to the answer?

OpenStudy (anonymous):

I agree to what you just said precal

terenzreignz (terenzreignz):

Agreed :) Allow me... \[\large S=a+ar+ar^2+ar^3+ar^4...\]\[\large \frac{S}r=\frac{a}r+a+ar+ar^3+ar^4+ar^5...\]\[\large \frac{S}r-\frac{a}r=a+ar+ar^2+ar^3+ar^4...\] And our usual result, the right side is once again equal to S :D \[\large \frac1r(S-a)=S\]\[\large S-a=Sr\]\[\large S-Sr=a\]\[\large S(1-r)=a\]\[\huge \color{blue}{S=\frac{a}{1-r}}\]

terenzreignz (terenzreignz):

So, plug in and be done with it,@Chad123 :D

terenzreignz (terenzreignz):

a is the first term in the sequence(series) and r is the common ratio

terenzreignz (terenzreignz):

a = 1/3 ?

OpenStudy (anonymous):

Sounds logical to me now

terenzreignz (terenzreignz):

What is the 3/4 doing there?

OpenStudy (anonymous):

100 % sure

terenzreignz (terenzreignz):

LOL \[\huge \frac{a}{\frac23}=\frac12 \]\[\huge \frac23\times \frac{a}{\frac23}=\frac23\times\frac12 \]\[\huge \cancel{\frac23}\times \frac{a}{\cancel{\frac23}}=\frac{\cancel{2}}{3}\times\frac{1}{\cancel{2}} \]

terenzreignz (terenzreignz):

Don't get dizzy, @precal XD

terenzreignz (terenzreignz):

a is 1/3 ?

terenzreignz (terenzreignz):

Your mistake this time was multiplying 3/2 to both sides instead of 2/3 which is the correct factor to cancel out the denominator 2/3

terenzreignz (terenzreignz):

That only applies when n starts at 0.

terenzreignz (terenzreignz):

a is the first term... don't get bogged down into shortcuts :D

terenzreignz (terenzreignz):